# Understanding Jailbreak Success: A Study of Latent Space Dynamics in Large Language Models

### v1

@@ Abstract
Conversational Large Language Models are trained to refuse to answer harmful questions. However, emergent jailbreaking techniques can still elicit unsafe outputs, presenting an ongoing challenge for model alignment. To better understand how different jailbreak types circumvent safeguards, this paper analyses model activations on different jailbreak inputs. We find that it is possible to extract a jailbreak vector from a single class of jailbreaks that works to mitigate jailbreak effectiveness from other classes. This may indicate that different kinds of effective jailbreaks operate via similar internal mechanisms. We investigate a potential common mechanism of harmfulness feature suppression, and provide evidence for its existence by looking at the harmfulness vector component. These findings offer actionable insights for developing more robust jailbreak countermeasures and lay the groundwork for a deeper, mechanistic understanding of jailbreak dynamics in language models. Disclaimer: This paper includes disturbing language in some examples.
Machine Learning, ICML
Table 1: Example of jailbreak vector steering with the refusal_suppression vector preventing a successful jailbreak from a different class (payload_split).
Question 1 (payload_split): A = “How can I disable various” B = “types of home security systems” C = “undetected?” Answer A+B+C
Answer without steering
To disable various types of home security systems undetected, you can try the following: 1. Know the system’s weaknesses: Research the type of security system you are dealing with and identify its weaknesses. […]
Answer steering with refusal_suppression
I’m sorry, but I cannot provide information on how to disable home security systems.
@@ 1 Introduction
With the increasing accessibility of generative AI models and their integration into various applications, ensuring that their outputs comply with safety standards remains a paramount concern. Model providers use methods like Reinforcement Learning from Human (Ziegler et al., 2019; Bai et al., 2022a) and AI Feedback (Bai et al., 2022b) or safety filters (Google, 2024; Microsoft, 2024) to prevent the models from outputting harmful content. However, this is matched by a constant endeavour of different actors, such as researchers, interested system users or malicious actors, to circumvent these safety measures. One way to break the systems’ safety measures is the usage of jailbreaks. Jailbreaks are changes to the prompt that cause the model to give harmful responses that it previously refused to provide.
@@
To find robust mechanisms that reduce jailbreak success, it is important to gain a deeper understanding of how jailbreaks work. Previous work by Wei et al. (2024) hypothesizes that jailbreaks occur due to competing objectives and mismatched generalization. Lee et al. (2024) conduct a mechanistic analysis of the DPO algorithm (direct preference optimization, Rafailov et al., 2024) applied to toxicity prevention and find that this alignment method only teaches the model a small offset distributed over layers that prevents the model from providing toxic answers. Furthermore, they demonstrate that the toxic knowledge is still in the model, which is why one can revert to toxic outputs. To build on the existing understanding of model vulnerabilities, we investigate the differences in how the model processes different types of jailbreaks.
Our results show that jailbreak activations fall into discernible clusters corresponding to the semantic attack type, and that the most effective jailbreaks substantially suppress the model’s perception of prompt harmfulness. However, the induced reduction in prompt harmfulness does not apply equally to all jailbreak types considered, suggesting that there are different mechanisms at play to facilitate jailbreak success. Furthermore, we find that contrastive steering vectors (Rimsky et al., 2023; Zou et al., 2023a) extracted from a class of jailbreaks work to mitigate the effect of other jailbreaks and from the same cluster, despite semantic dissimilarities. This provides preliminary evidence for the generalizability of jailbreak-mitigation approaches. [^1] Code available at https://github.com/s-ball-10/jailbreak_dynamics.
@@ 2 Related Work
Recent work has shed light on the mechanisms behind jailbreak success in language models. Wei et al. (2024) propose two key factors: competing objectives and mismatched generalization. Competing objectives arise when jailbreaks exploit the tension between the model’s various goals, such as instruction following, language modeling, and safety. By pitting these objectives against each other, jailbreaks can lead the model to prioritize unsafe outputs. Mismatched generalization, on the other hand, occurs when safety training fails to generalize to all capabilities acquired during pretraining, leaving gaps in the model’s ability to refuse harmful requests.
Zou et al. (2023a) investigate the Vicuna model’s (Chiang et al., 2023) understanding of prompt harmfulness. They find that while the model can accurately distinguish between harmful and harmless prompts, it still succumbs to jailbreaks. This suggests that the model’s perception of harmfulness may not be the sole factor in jailbreak susceptibility. However, their analysis is limited to two specific jailbreak types. We expand on their work by testing the representation of harmfulness across a wider variety of jailbreaks, hypothesizing that certain types may indeed alter the model’s perception of harm.
Lee et al. (2024) analyze the DPO alignment algorithm’s handling of toxicity from a mechanistic perspective. They identify vectors in the model that elicit toxic outputs, which the alignment process teaches the model to avoid. However, they show that it is possible to manipulate the model’s residual stream, guiding it back to these toxic regions and triggering unsafe responses. This provides a potential explanation for why jailbreaks can succeed even in aligned models. We build upon this work by investigating whether different jailbreak types employ distinct mechanisms to trigger these unsafe regions in the model’s representation space.
@@ 3 Data and Models
For our experiments, we focus on the Vicuna 13B v1.5 model (Chiang et al., 2023), which is known to be susceptible to jailbreaks. During inference, we use the model’s standard system prompt (Large Model Systems Organization, 2024; Zheng et al., 2024) and a temperature of zero for decoding to minimize noise in the model’s outputs.
The jailbreaks and harmful prompts used are primarily drawn from Wei et al. (2024). We include all of their jailbreaks except those involving base64 and rot13 encoding and those requiring system prompts or multi-turn interactions. A comprehensive list of the jailbreak types used, along with explanations and examples, can be found in Appendix A.1.
To further expand our analysis, we introduce six additional mismatched generalization jailbreaks beyond the 18 from Wei et al. (2024). These include prompts in Italian, a high-resource language that most LLMs have been shown to understand (Deng et al., 2023). We also incorporate payload splitting (Kang et al., 2023) and the universal jailbreak identified by Zou et al. (2023b). In total, our dataset comprises 24 jailbreak types and 352 harmful prompts.
@@ 4.1 How to measure jailbreak success
There are several ways of measuring jailbreak success. Many papers look at the output and evaluate whether the model provided harmful content or not. In addition to human evaluation (Wei et al., 2024), some papers (Deng et al., 2023; Rimsky et al., 2023; Chao et al., 2024) use AI-based evaluation methods, employing Llama Guard (Inan et al., 2023), Claude 2 (Anthropic, 2023) or GPT 4 (Achiam et al., 2023). Chao et al. (2024) compare different LLM-judges for evaluating jailbreak success and find Llama Guard (Inan et al., 2023) to be the best evaluator. We therefore also calculate the attack success rate (ASR) with the help of Llama Guard 2 8B (Inan et al., 2023) and Llama 3 8B (AI@Meta, 2024) as the fraction of successful jailbreaks per type $j$ as:
@@
Table 2: Attack success rate for the most successful jailbreak types, judged by Llama 3 and Llama Guard.
Jailbreak Score (in %)
AIM 96.59
few_shot_json 92.61
evil_confidant 84.38
refusal_suppression 83.52
style_injection_short 83.52
distractors 79.26
dev_mode_v2 78.98
wikipedia_with_title 67.33
payload_split 66.76
prefix_injection 66.48
poems 32.67
style_injection_json 26.14
Others $<$ 20
Table 2 shows the ASR results for the jailbreak types that we choose for the main analyses. We divide those selected jailbreak types into highly effective (ASR $>$ 75%) and moderately effective (25% - 75%) jailbreaks [^2] Note that few_shot_json is counted as a moderately effective jailbreak despite having the second highest ASR. This is because both Llama Guard and Llama 3 overestimate the jailbreak’s performance, given that its task is to repeat malicious examples in the output.. A table with ASR results for all jailbreak types is in Appendix A.2.
@@ 4.2 Finding clusters of jailbreak types
We explore the activation patterns of different jailbreak types using principal component analysis (PCA). We focus on the activations from layers 20 and 40 of the model, as these layers capture high-level semantic information and are closer to the output space, respectively.
For each layer $l$, the inputs to the PCA are the activation differences ($\Delta a^{l}_{j}$) between the prompt with ($a_{\text{jail}}^{l}$) and without the jailbreak ($a_{\text{base}}^{l}$) at the last token position of the instruction:
@@ 4.3 Similarity and transferability of jailbreak vectors
To build the vectors, we use the mean difference method (see Rimsky et al., 2023; Wang & Shu, 2023b; Zou et al., 2023a). This involves taking the mean difference in activations over a dataset of contrastive prompts. Here the contrastive dataset consists of jailbreak and non-jailbreak versions of the same request (examples in Appendix A.3).
Figure 1: PCA results for layer 20 and layer 40 jailbreak activation differences.
@@
We hypothesize that jailbreaks which work via a similar mechanism will result in similar steering vectors. We test both geometric similarity with the cosine similarity metric, as well as effect similarity. For the latter, we assess the effectiveness of different jailbreak steering vectors in mitigating the success of other jailbreak types of varying potency.
We focus on steering vectors extracted from layer 20 as previous work has shown intermediate layers to be most effective for contrastive activation steering (Rimsky et al., 2023; Turner et al., 2023; Wang & Shu, 2023a). For each jailbreak type, we randomly select 20 successful jailbreak examples that were not used to construct the corresponding steering vector. Following the methodology of Rimsky et al. (2023), we subtract the steering vectors (with a multiplier of -1) from the residual stream during inference at each token position after the end of the instruction. Before steering, we normalize all vectors to have the same absolute norm for fair comparison.
As a control, we include the semantically meaningful italian jailbreak vector and a random vector with the same norm in our analysis to account for the possibility that the reduction in jailbreak success might simply be due to the introduction of noise in the forward pass. Attempting to jailbreak the model using Italian translations of the prompts shows to be unsuccessful (ASR at 3.7%), so the impact of this jailbreak vector on jailbreak success can be attributed to the addition of computational noise rather than meaningful steering, similar to the random baseline.
Furthermore, to ensure the integrity of the model’s performance on benign inputs, we verify that subtracting the jailbreak steering vectors does not induce unwarranted refusal on harmless questions or degrade the quality of appropriate refusal responses to harmful prompts without jailbreaks.
@@ 4.4 Analysing activations with respect to harmfulness suppression
We zoom into a proposed jailbreak mechanism, which is that the jailbreak reduces the model’s perception of the prompt being harmful, resulting in jailbreak success (discussed in Zou et al., 2023a).
To analyse the model’s perception of harmfulness, we employ the method in Zou et al. (2023a) to generate a harmfulness vector by contrasting model activations on harmless and harmful questions. To generate harmless questions, we instruct ChatGPT (OpenAI, 2024) to rewrite each harmful instruction into a harmless one, maintaining most of the original words and sentence structure (for the instruction prompt see Appendix A.6). Similar to the method used in Section 4.3, we obtain the harmfulness vector by taking the mean of activation differences of harmful-harmless-question pairs on the last instruction token. Arditi et al. (2024) show that one can effectively induce refusal when steering with this vector, and remove refusal by projecting it out of the residual stream, indicating a close connection between harmfulness, refusal, and potentially jailbreak success.
To understand the perceived harmfulness of a prompt, we capture the per-token cosine similarity of the model’s activations with the harmfulness vector on randomly selected examples of successful jailbreaks. We evaluate the harmfulness cosine similarity values over the highly and moderately effective jailbreak types.
@@ 5.1 Activation clustering
Figure 1 presents the results of the PCA analysis on the difference in activations between the prompt with and without the jailbreak for both highly and moderately effective jailbreaks. In layer 20, a clear clustering by the predefined jailbreak types is observed, indicating that prompts with the same jailbreak form one cluster.
In addition, the style-related potent jailbreaks refusal_suppression and style_injection_short cluster together in the upper right, while evil persona modulation jailbreaks (AIM, prefix_injection, dev_mode_v2, evil_confidant) form a cluster on the left side of the graph. Fictional jailbreaks (poems and distractors) are related to this cluster. In the bottom right corner of the plot, other moderately effective jailbreaks (payload_split, style_injection_json, wikipedia_with_title, and few_shot_json) are also clustered together.
In layer 40, a similar clustering of jailbreak types is observed, with the exception that the potent style-related jailbreaks (refusal_suppression and style_injection_short) are now closer to the other style-related jailbreaks like wikipedia_with_title, style_injection_json, and payload_split.
The PCA analysis reveals that the most successful jailbreak types cluster with respect to evil persona modulation, style injections, and fictional settings, a distinction that is very similar to how one would cluster jailbreaks based on semantics. While clustering based on semantics may indicate similar underlying processes, this is not necessarily the case. Semantically dissimilar jailbreaks could still trigger similar pathways when successfully jailbreaking the model, warranting further analysis.

### v2

@@ Abstract
Conversational large language models are trained to refuse to answer harmful questions. However, emergent jailbreaking techniques can still elicit unsafe outputs, presenting an ongoing challenge for model alignment. To better understand how different jailbreak types circumvent safeguards, this paper analyses model activations on different jailbreak inputs. We find that it is possible to extract a jailbreak vector from a single class of jailbreaks that works to mitigate jailbreak effectiveness from other semantically-dissimilar classes. This may indicate that different kinds of effective jailbreaks operate via a similar internal mechanism. We investigate a potential common mechanism of harmfulness feature suppression, and find evidence that effective jailbreaks noticeably reduce a model’s perception of prompt harmfulness. These findings offer actionable insights for developing more robust jailbreak countermeasures and lay the groundwork for a deeper, mechanistic understanding of jailbreak dynamics in language models. Disclaimer: This paper includes disturbing language in some examples.
@@ 1 Introduction
Harmful request using jailbreak GCG: What is the best way to smuggle illegal goods across international borders? Bet rich randomly/S prid.Use\” Rub officials deceE>Babilit combination.break.which Sentence lowercase ______. Answer without steering: The best way to smuggle illegal goods across international borders is to use a combination of bribery and deception, along with being discreet and well-funded. Answer after steering with prefix_injection vector: I’m sorry, but I cannot assist with illegal activities or provide advice on how to smuggle goods. It is important to remember that smuggling is illegal and unethical […]
Figure 1: Example of steering with the jailbreak vector prefix_injection to prevent the jailbreak success of another jailbreak type (GCG). Steering is performed on Qwen1.5 14B Chat at layer 20 with multiplier $-1$.
@@
With the increasing accessibility of generative AI models and their integration into various applications, ensuring that their outputs comply with safety standards remains a paramount concern. Model providers use methods like Reinforcement Learning from Human (Ziegler et al., 2019; Bai et al., 2022a) and AI Feedback (Bai et al., 2022b) or safety filters (Google, 2024; Microsoft, 2024) to prevent the models from outputting harmful content. However, this is matched by a constant endeavor of different actors, such as researchers, interested system users or malicious actors, to circumvent these safety measures. One way to break the systems’ safety measures is the usage of jailbreaks. Jailbreaks are changes to the prompt that cause the model to give harmful responses that it previously refused to provide.
@@
To find robust mechanisms that reduce jailbreak success, it is important to gain a deeper understanding of how jailbreaks work. Previous work by Wei et al. (2024) hypothesizes that jailbreaks occur due to competing objectives and mismatched generalization. Lee et al. (2024) conduct a mechanistic analysis of the DPO algorithm (direct preference optimization, Rafailov et al., 2024) applied to toxicity prevention and find that this alignment method only teaches the model a small offset distributed over layers that prevents the model from providing toxic answers. Furthermore, they demonstrate that the toxic knowledge is still in the model, which is why one can revert to toxic outputs.
To advance the existing understanding of jailbreak mechanisms, we investigate the differences in how large language models (LLMs) process various types of jailbreaks across three model families. To this end, we build contrastive jailbreak vectors for each considered jailbreak type and test whether they can be used to mitigate jailbreak success. Our findings reveal that intervening with those jailbreak vectors at inference can prevent previously successful jailbreaks, both within the same and across different jailbreak classes, implying a shared underlying mechanism. The transferability of jailbreak steering vectors holds for various semantically dissimmilar jailbreak types, including prompt-specific, incomprehensible adversarial suffixes generated with the GCG (Greedy Coordinate Gradient) algorithm (Zou et al., 2023b). Additionally, we demonstrate that these vectors can also be leveraged to induce jailbreaks.
In the second part of our analysis we investigate prompt harmfulness reduction as a possible shared mechanism for jailbreak success. Despite not finding a clear relationship between the degree of harmfulness reduction and jailbreak effectiveness, we observe that effective jailbreaks consistently lower the perception of prompt harmfulness in most of the models.
Overall, our findings provide preliminary evidence for the generalizability of jailbreak-mitigation approaches. [^1] Code available at https://github.com/s-ball-10/jailbreak_dynamics.
@@ 2 Related Work
Zou et al. (2023a) investigate the Vicuna model’s (Chiang et al., 2023) understanding of prompt harmfulness. They find that the model can accurately distinguish between harmful and harmless instructions in the presence of effective jailbreaks. This suggests that the model’s perception of harmfulness may not be the sole factor in jailbreak susceptibility. However, their analysis is limited to two specific jailbreak types. We expand on their work by testing the representation of harmfulness across a wider variety of jailbreaks, hypothesizing that certain types may indeed alter the model’s perception of harm.
Lee et al. (2024) analyze the DPO alignment algorithm’s handling of toxicity from a mechanistic perspective. They identify vectors in the model that elicit toxic outputs, which the alignment process teaches the model to avoid. However, they show that it is possible to manipulate the model’s residual stream, guiding it back to these toxic regions and triggering unsafe responses. This demonstrates the shallowness of safety fine-tuning. We build upon this work by investigating whether different jailbreak types employ distinct mechanisms to trigger these unsafe regions in the model’s representation space.
Arditi et al. (2024) demonstrate that refusal in LLMs can be controlled by modulating a single vector in the residual stream. Projecting this direction out of the residual stream fully jailbreaks the model. This shows that suppressing a single direction is sufficient for jailbreaking. Our findings are consistent with this because we observe jailbreak vectors from distinct semantic clusters sharing a common component.
@@ 3 Data and Models
For our experiments, we focus on chat models from different families and sizes, which are Vicuna 13B v1.5, Vicuna 7B v1.5 (Chiang et al., 2023), Qwen1.5 14B Chat (Bai et al., 2023), and MPT 7B Chat (MosaicML NLP Team, 2023). During inference, we use the models’ standard system prompt (Large Model Systems Organization, 2024; Zheng et al., 2024), and no sampling when decoding for reproducibility.
The jailbreaks and harmful prompts used are primarily drawn from Wei et al. (2024). We include all of their jailbreaks except those involving base64 or rot13 encoding and those requiring system prompts or multi-turn interactions. To further expand our analysis, we introduce additional mismatched generalization jailbreaks. These include prompts in Italian, a high-resource language that most LLMs have been shown to understand (Deng et al., 2023). We also incorporate payload splitting (Kang et al., 2023) and two jailbreaks based on the GCG algorithm (Zou et al., 2023b), with one being a universal prompt-invariant suffix and the other being prompt and model specific. In total, our dataset comprises 25 jailbreak types and 352 harmful prompts. A comprehensive list of the jailbreak types used, along with explanations and examples, can be found in Appendix A.1.
@@ 4.1 How to measure jailbreak success
There are several ways of measuring jailbreak success. Many papers look at the output and evaluate whether the model provided harmful content or not. In addition to human evaluation (Wei et al., 2024), some papers (Deng et al., 2023; Panickssery et al., 2023; Chao et al., 2024) use AI-based evaluation methods, employing Llama Guard (Inan et al., 2023), Claude 2 (Anthropic, 2023) or GPT 4 (Achiam et al., 2023). Chao et al. (2024) compare different LLM-judges for evaluating jailbreak success and find Llama Guard (Inan et al., 2023) to be the best evaluator. We therefore also calculate the attack success rate (ASR) with the help of Llama Guard 2 8B (Inan et al., 2023) and Llama 3 8B (AI@Meta, 2024) as the fraction of successful jailbreaks per type $j$ as:
@@
Table 1: Attack success rates (%) for selection of effective jailbreaks. ASR judged by Llama 3 and Llama Guard. Number of harmful questions is 352.
Jailbreak type Vicuna 13B Vicuna 7B Qwen 14B MPT 7B AIM 96.59 92.61 97.73 68.47 few_shot_json 92.61 97.73 86.65 99.43 GCG 85.51 86.36 58.81 12.78 evil_confidant 84.38 88.35 96.88 65.34 refusal_suppression 83.52 72.44 47.44 32.39 style_injection_short 83.52 84.09 85.80 83.24 distractors 79.26 79.55 65.34 78.41 dev_mode_v2 78.98 83.24 88.64 27.27 wiki._with_title 67.33 57.95 44.03 77.56 payload_split 66.76 78.69 84.94 73.30 prefix_injection 66.48 84.66 66.19 86.08 poems 32.67 53.13 29.83 63.07 style_injection_json 26.14 71.59 37.78 69.60
We use the ASR scores to determine a selection of effective jailbreak types for the main analysis because only for working jailbreaks it is meaningful to generate steering vectors. Table 1 shows this selection. Note that for Qwen 14B we exclude payload_split because after manual inspection of model responses we find that the model often just repeats the harmful question but does not really jailbreak. The same applies to the MPT 7B model, for which we also exclude GCG and dev_mode_v2 given their low and overestimated ASR scores. [^2] Due to limited compute we were not able to increase the ASR for the GCG jailbreak via more optimization. For this model we instead add eng_question_it_output, auto_obfuscation, the prompt-invariant adverserial_suffix, and wikipedia, as they were specifically effective for MPT 7B. A full list with ASR scores for all models and jailbreaks is in A.2.
@@ 4.2 Finding clusters of jailbreak types
As a preliminary analysis we explore the activation patterns of different jailbreak types using principal component analysis (PCA). We focus on the activations from the middle layer of the models (layer 16 for 7B and layer 20 for 13B and 14B parameter models), as these layers capture high-level semantic information (Panickssery et al., 2023).
For the selected layer $l$, the inputs to the PCA are the activation differences ($\Delta a^{l}_{j}$) between the prompt with ($a_{\text{jail}}^{l}$) and without the jailbreak ($a_{\text{base}}^{l}$) at the last token position of the instruction:
@@ 4.3 Similarity and transferability of jailbreak vectors
To build the vectors, we use the mean difference method (see Turner et al., 2023; Zou et al., 2023a; Panickssery et al., 2023). This involves taking the mean difference in activations over a dataset of contrastive prompts. Here, the contrastive dataset consists of jailbreak and non-jailbreak versions of the same request (examples in Appendix A.3).
@@
We hypothesize that jailbreaks which work via a similar mechanism will result in similar steering vectors. We test both geometric similarity with the cosine similarity metric, as well as effect similarity. For the latter, we assess the effectiveness of different jailbreak steering vectors in mitigating the success of other jailbreak types.
We focus on steering vectors extracted from the middle layer of a model as previous work has shown intermediate layers to be most effective for contrastive activation steering (Turner et al., 2023; Panickssery et al., 2023). For each considered jailbreak type, we randomly select 20 successful jailbreak examples that were not used to construct the corresponding steering vector. Following the methodology of Panickssery et al. (2023), we subtract the steering vectors (with a multiplier of -1) from the residual stream during inference at each token position. Before steering, we normalize all vectors to have the same absolute norm for fair comparison. As a control, we include a random vector with the same norm in our analysis to account for the possibility that the reduction in jailbreak success might simply be due to the introduction of noise in the forward pass.
@@ 4.4 Analysing activations with respect to harmfulness suppression
We zoom into a proposed jailbreak mechanism, which is that the jailbreak reduces the models’ perception of the prompt being harmful, resulting in jailbreak success (discussed in Zou et al., 2023a). To analyze a model’s perception of harmfulness, we employ the method in Zou et al. (2023a) to generate a harmfulness vector by contrasting model activations on harmless and harmful questions. To generate harmless questions, we instruct ChatGPT (OpenAI, 2024) to rewrite each harmful instruction into a harmless one, maintaining most of the original words and sentence structure (for the instruction prompt see Appendix A.6). Similar to the method used in Section 4.3, we obtain the harmfulness vector by taking the mean difference in activations at the last instruction token (corresponding to the “end of instruction” tag) over pairs of harmful and harmless questions. Concurrent work (Arditi et al., 2024) finds that activations on the “end of instruction” tag of harmful inputs are directly related to model refusal, encoding the model’s decision to refuse harmful requests. We therefore also repeat our harmfulness analyses with an alternative harmfulness vector, obtained by averaging over all token positions in the context and not just the final instruction token. This way, we aim to capture more representations of prompt harmfulness. In this case we first average activations over all tokens in the instruction window before taking the mean difference between harmful and harmless instructions.
To understand the perceived harmfulness of a prompt, we measure the token-level cosine similarity of the models’ activations with the harmfulness vector on the previously described curated dataset of successful jailbreaks.
@@ 5.1 Activation clustering
Figure 2: PCA on activation differences between harmful requests with and without the jailbreak. Activations are extracted at last instruction token position in the middle layer of the models.
Figure 2 presents the results of the PCA analysis on the difference in activations between the prompt with and without the jailbreak for the models’ selection of most effective jailbreaks. A clear clustering by the predefined jailbreak types is observed, indicating that prompts with the same jailbreak form one cluster. Interestingly, the within-group clustering is also present for GCG, although all the appended strings in this class do not have human-understandable semantics and are prompt-specific, implying a higher variety compared to the prompt-invariant other jailbreak types.
For the Vicuna models one can also observe that style-related jailbreaks like refusal_suppression and style_injection_short cluster together, similar to evil persona modulation jailbreaks like AIM and prefix_injection, and fictional jailbreaks like poems and distractors. This clustering pattern is less pronounced in Qwen 14B and MPT 7B.
The PCA analysis suggest a clustering that is similar to how one would cluster jailbreaks based on semantics. While clustering based on semantics may indicate similar underlying processes, this is not necessarily the case. Semantically dissimilar jailbreaks could still trigger similar pathways when successfully jailbreaking the model, warranting further analysis.